OpenStudy (anonymous):

1. Write the equation of the line with a slope of 2/3 and a y-intercept of 5 in general form. Clearly show and explain each step

OpenStudy (anonymous):

@jazy ? can u help

OpenStudy (anonymous):

given m=2/3 and c=5 hence both slope and y -intercpt of y=mx +c is known ths the required eq of line is y=2/3 x +5 or y-5 =2x/3 or 3y-15=2x or 2x -3y +15=0 (the required eq. of line )

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

So you need it in standard form?

OpenStudy (anonymous):

Or slope-intercept form?

OpenStudy (anonymous):

slope intercept

OpenStudy (anonymous):

Anyways we need to get it into slope-intercept before any other format. So you have a slope of 2/3 and a y-intercept of 5. Remember that Slope-intercept form is y = mx + b where m is slope and b is y-intercept. So just plug in the values. y = 2/3x + 5

OpenStudy (anonymous):

so thats the answer? all i do is plug it in?

OpenStudy (anonymous):

Pretty much... If its Slope-intercept form what you're looking for.

OpenStudy (anonymous):

thank you